# Automatic Tracing in Task-Based Runtime Systems

### v1

@@ 6.3. Overheads of Apophenia
To measure the time taken until Apophenia reaches a steady state of replaying traces on our iterative applications, we report the number of iterations until a steady state is reached. Figure 9 contains the iteration counts needed for each application in Section 6.1 and Section 6.2, which range from 30 to 300. These simulation and machine learning workloads would be run in production for a significantly larger number of iterations. We note that the cuNumeric applications have a larger number of required warmup iterations due to the dynamic behavior discussed in Section 2, where a single application-level iteration of the program does not necessarily correspond to a repeated sequence of tasks.
@@ 7. Related Work
Just-In-Time (JIT) compilers (Hölzle and Ungar, 1996; Gal et al., 2009; Paleczny et al., 2001) for dynamic languages have a tiered execution system, where the target language is first translated to bytecode, which is executed by an interpreter. Frequently executed program fragments are then compiled into native instructions for significantly faster execution. Apophenia employs a similar architecture where a task-based runtime system’s dynamic analysis acts as the slow but general interpreter, and uses a tracing engine as the fast but specialized compiler. JIT compilers rely on code landmarks like function definitions and basic block addresses to maintain counters of frequently executed program fragments. Since Apophenia views an unrolled stream of tasks, it must employ novel techniques for identification of traceable program fragments.
@@
Several task-based runtime systems have been developed for high performance computing (Bauer et al., 2012; Augonnet et al., 2011; Bosilca et al., 2011), data science (Dask Development Team, 2016; Zaharia et al., 2012), and machine learning (Moritz et al., 2017; Barham et al., 2022). One axis of runtime overhead that these different systems impose on applications is the cost of dependence analysis. The cost of dependence analysis is directly related to the expressivity and flexibility of the runtime system’s programming model. Legion has an expressive data model that supports content-based coherence (Bauer et al., 2023), leading to a relatively expensive dependence analysis. As a result, tracing (Lee et al., 2018) was developed to reduce the costs of the dependence analysis. A tracing-like technique called Execution Templates (Mashayekhi et al., 2017) was also developed to cache control plane decisions in runtime systems for cloud-based environments.

### v2

@@ 6.3. Overheads of Apophenia
To measure the time taken until Apophenia reaches a steady state of replaying traces on our iterative applications, we report the number of iterations until a steady state is reached. Figure 9 contains the iteration counts needed for each application in Section 6.1 and Section 6.2, which range from 30 to 300. These simulation and machine learning workloads would be run in production for a significantly larger number of iterations, so speedup in the steady state corresponds closely to end-to-end speedup. We note that the cuPyNumeric applications have a larger number of required warmup iterations due to the dynamic behavior discussed in Section 2, where a single application-level iteration of the program does not necessarily correspond to a repeated sequence of tasks.
In terms of resource utilization, Apophenia requires a modest amount of CPU memory to store the history buffer of tasks for analysis. Apophenia runs the asynchronous string analysis (Section 4.2) on Legion’s background worker threads. We have not found these resource requirements to impact application performance or memory utilization.
@@ 7. Related Work
Just-In-Time (JIT) compilers (Hölzle and Ungar, 1996; Gal et al., 2009a; Paleczny et al., 2001) for dynamic languages have a tiered execution system, where the target language is first translated to bytecode, which is executed by an interpreter. Frequently executed program fragments are then compiled into native instructions for significantly faster execution. Apophenia employs a similar architecture where a task-based runtime system’s dynamic analysis acts as the slow but general interpreter, and uses a tracing engine as the fast but specialized compiler.
Tracing-based JIT compilers such as TraceMonkey (Gal et al., 2009b) record sequences of instructions executed at runtime and generate optimized code for those sequences. Method-based JIT compilers identify frequently invoked functions in the target program and compile type-specialized versions of those functions. JIT compilers identify the desired instruction sequences or methods to compile by relying on code landmarks like function definitions and basic block addresses to maintain counters of frequently executed program fragments. Since Apophenia views an unrolled stream of tasks, it must employ novel techniques for identification of traceable program fragments.
JIT compilers also perform dynamic analysis to recover data structures like call-graphs from the target program. Sampling-based methods (Zhuang et al., 2006) have been developed to balance runtime cost of profiling each function call with the accuracy of the sampled data structure. Discovering traces in our work requires for long contiguous sequences of issued tasks to be analyzed together, as a trace must repeat several times to be considered by Apophenia. Breaking up these sequences with independent and non-contiguous samples can lead to a loss of precision when discovering traces. Instead, Apophenia employs an always-on approach where all tasks are analyzed, and uses a sub-sampling method on the set of collected tasks to manage the trade off between responsiveness of the trace analysis and length of the discovered traces. An always-on approach is cheaper to use in the task-based runtime system context than within a standard JIT compiler as tasks are relatively coarse when compared to bytecode instructions.
@@
Several task-based runtime systems have been developed for high performance computing (Bauer et al., 2012; Augonnet et al., 2011; Bosilca et al., 2011), data science (Dask Development Team, 2016; Zaharia et al., 2012), and machine learning (Moritz et al., 2017; Barham et al., 2022). One axis of runtime overhead that these different systems impose on applications is the cost of dependence analysis. The cost of dependence analysis is directly related to the expressivity and flexibility of the runtime system’s programming model. Legion has an expressive data model that supports content-based coherence (Bauer et al., 2023), leading to a relatively expensive dependence analysis. As a result, tracing (Lee et al., 2018) was developed to reduce the costs of the dependence analysis. Both the StarPU (Agullo et al., 2017) and PARSEC (Hoque et al., 2017) runtime systems have modes that perform a dynamic dependence analysis to extract parallelism, and these modes have been shown to add overheads over the explicitly-parallel, analysis free modes (Slaughter et al., 2020). Tracing techniques could be applied within these runtime systems to lower the overheads of the dynamic, implicitly parallel modes.
Techniques similar to tracing have also been developed in other runtime systems to lower overheads. A tracing-like technique called Execution Templates (Mashayekhi et al., 2017) was developed to cache control plane decisions in runtime systems for cloud-based environments. The Dask (Dask Development Team, 2016) runtime system exposes an API for users to explicitly construct and optimize task graphs (Dask, 2024), which is lower-level but more efficient than the standard individual task launching API. The Ray (Moritz et al., 2017) runtime system has recently added an execution mode called “Compiled Graphs” (Authors, 2024), where users build explicit computation graphs and issue them to Ray for lower overhead replay. Finally, the CUDA runtime exposes a similar feature to tracing called CUDA Graphs (NVIDIA, 2024), where users may record a sequence of CUDA kernel launches and replay the sequence with lower overheads. Techniques used in Apophenia could potentially be applied to these systems to remove the requirement for users to be involved in the memoization and optimization of these computations.
@@ 8. Conclusion
Acknowledgements
We thank Wonchan Lee, Manolis Papadakis and Shriram Jagannathan for their assistance with Legate. We thank Seshu Yamajala for his assistance in running the S3D simulation. We thank Elliot Slaughter for his assistance in debugging and running Regent programs. We thank Mario Di Renzo and Caetano Melone for their assistance in running the HTR simulation. We thank Zhihao Jia and Colin Unger for their assistance in running FlexFlow. We thank Roshni Sahoo for her assistance in developing formal optimization problem in Section 3. We thank Danny Sleator and Sam Westrick for suggestions and pointers to related work around the string analysis component of this work. We thank Wei Wu for discussions about the PARSEC runtime system, and Cedric Augonnet for discussions about the StarPU runtime system. We thank (in no particular order) James Dong, AJ Root, Chris Gyurgyik, Rubens Lacouture, Shiv Sundram, Scott Kovach and Olivia Hsu for their discussions and feedback on this manuscript. Rohan Yadav was supported by an NVIDIA Graduate Fellowship, and part of this work was done while Rohan Yadav was an intern at NVIDIA Research. This work was in part supported by the National Science Foundation under Grant CCF-2216964 and by Digital Futures at KTH Royal Institute of Technology. This research used resources of the National Energy Research Scientific Computing Center, a DOE Office of Science User Facility supported by the Office of Science of the U.S. Department of Energy under Contract No. DE-AC02-05CH11231 using NERSC award ASCR-ERCAP0026353.
Appendix A Artifact Appendix
A.1. Abstract
This artifact presents the computational artifact of Apophenia, a system that automatically traces Legion applications. This artifact is supported by an implementation of Apophenia within the Legion runtime system and a standalone implementation of the repeated sub-strings algorithm as described in the paper. We also provide artifacts for the subset of our benchmarks that are open-source.
We evaluated Apophenia on the Eos and Perlmutter supercomputers. Each node of Eos is an NVIDIA DGX H100, containing 8 H100 GPUs with 80 GB of memory and a 112 core Intel Xeon Platinum. Each node of Perlmutter contains 4 NVIDIA A100 GPUs with 40 GB of memory and a 64 core AMD EPYC 7763. Nodes of Eos are connected with an Infiniband interconnect, while Perlmutter uses a Slingshot interconnect. We compile Legion on Eos with the UCX networking module, and use the GASNet-EX networking module on Perlmutter.
A.2. Artifact check-list (meta-information)
A.3. Description
A.3.1. How to access
The version of Legion with Apophenia is available here. The standalone implementation of the repeated substrings algorithm is available here. The version of TorchSWE used for benchmarking is available here, though executing it with Apophenia requires a currently closed-source version of the Legate runtime and cuNumeric. The version of FlexFlow used for benchmarking is available here.
A.3.2. Hardware dependencies
Our experiments were run on server-class machines with multiple GPUs per node. While these are not necessary, the scaling and problem sizes that fit on each node will differ on different setups.
A.3.3. Software dependencies
We run all our experiments with Python 3.11. Aside from that, a standard supercomputer software stack (C++ compiler, CUDA installation, MPI installation) is expected.
A.4. Installation
Apophenia can be built and run with a standard Legion build, using the version of Legion from the artifact. The exact parameters to build Legion depend on the machine configuration. A sample Legion build for an Infiniband-based cluster with NVIDIA GPUs would invoke:
cd Legion/language
USE_CUDA=1 CONDUIT=ibv ./scripts/setup_env.sh
Since FlexFlow is fully open-source, we also provide build instructions. Using the provided version of Legion, FlexFlow can be built and installed for an NVIDIA machine with
export FF_GPU_BACKEND="cuda"
conda create -n flexflow
source activate flexflow
conda install -c conda-forge cmake make pillow \
  cmake-build-extension pybind11 numpy pandas \
  keras-preprocessing onnx transformers>=4.31.0 \
  sentencepiece einops
conda install -c pytorch pytorch torchvision torchaudio
conda install rust
pip3 install tensorflow notebook
cd FlexFlow
mkdir build
cd build
../config/config.linux
make -j
A.5. Experiment workflow
As a majority of our experiments are closed source, we do not provide a script that can run the full experiment suite. We do provide a command line that can be used to run the FlexFlow benchmark. The given command line is intended for SLURM based clusters, but additional configuration may be required depending on SLURM setup.
srun -N <NODES> \
  FlexFlow/build/examples/cpp/candle_uno/candle_uno \
  --warmup 30 \
  --batch-size 16384 \
  -ll:gpu <GPUS-PER-NODE> \
  -ll:fsize <GPU-MEM-IN-MBS> \
  -ll:util 4 \
  -ll:csize 30000 \
  -ll:zsize 5000 \
  -lg:enable_automatic_tracing \
  -lg:auto_trace:min_trace_length 25 \
  -lg:auto_trace:max_trace_length 200 \
  -lg:auto_trace:batchsize 5000 \
  -lg:auto_trace:identifier_algorithm \
      multi-scale \
  -lg:auto_trace:multi_scale_factor 500 \
  -lg:auto_trace:repeats_algorithm \
      quick_matching_of_substrings \
  -lg:inline_transitive_reduction \
  -lg:window 30000
A.6. Evaluation and expected results
We expect that when used on our benchmark applications, Apophenia finds and replays traces. On problem sizes that are Legion runtime-limited, this should result in speedup.
A.7. Experiment customization
Apophenia exposes several runtime configurations that are accepted by every Legion application for customizing the behavior.
  1. (1)
    -lg:enable_automatic_tracing: enable automatic tracing.
  2. (2)
    -lg:auto_trace:min_trace_length <N>: minimum length trace to consider.
  3. (3)
    -lg:auto_trace:max_trace_length <N>: maximum length trace to replay.
  4. (4)
    -lg:auto_trace:batchsize <N>: size of the task history buffer.
  5. (5)
    -lg:auto_trace:multi_scale_factor <N>: minimum size of the adaptive analysis.